A 20-minute Oil and Lube service claims that their average service time is no more than 20 minutes. A random sample of 40 service times was collected, and the sample mean was found to be 21.3 minutes, with a sample standard deviation of 5.5 minutes. Is there evidence to reject the claim at the alpha = 0.05 level? Perform an appropriate hypothesis test, showing the necessary calculations and/or explaining the process used to obtain the results.
Answer by stanbon(57361)

A 20-minute Oil and Lube service claims that their average service time is no more than 20 minutes.
---------------------------------
A random sample of 40 service times was collected, and the sample mean was found to be 21.3 minutes, with a sample standard deviation of 5.5 minutes.
-------------------------
Is there evidence to reject the claim at the alpha = 0.05 level?
Perform an appropriate hypothesis test, showing the necessary calculations and/or explaining the process used to obtain the results.
----------------------------
Ho: u <= 20 min (claim)
Ha: u > 20 min
--------------------
test statistic:
t(21.3) = (21.3-20)/[5.5/sqrt(40)] = 1.4949
------
p-value = P(t > 1.4949 when df=39) = 0.0715
---
Since the p-value is greater than 5%, fail to reject Ho.
Conclusion: The test results support the claim.
================
Cheers,
Stan H.